OpenStudy (natsent06):

PLZ HELP NEED ANSWER FAST WILL GIVE MEDAL suppose 30=a*60. So what is 30/a (30 over a)

OpenStudy (dangerousjesse):

30/.5

OpenStudy (dangerousjesse):

Or 15/1

OpenStudy (natsent06):

oh it is .5

OpenStudy (natsent06):

thanks I have been waiting for help forever

OpenStudy (anonymous):

does this really say \[30=a\times 60\]?

OpenStudy (dangerousjesse):

No problem. Do you know how I got this answer?

OpenStudy (natsent06):

satellite yes

OpenStudy (natsent06):

um not really

OpenStudy (anonymous):

if so then \(a=\frac{1}{2}\) and so \[\frac{30}{a}=\frac{30}{\frac{1}{2}}=30\times 2=60\]

OpenStudy (anonymous):

@DangerousJesse thirty divided by point five is sixty

OpenStudy (natsent06):

wait

OpenStudy (anonymous):

btw i loved a wrinkle in time mrs who was a crossword answer this week

OpenStudy (natsent06):

my thing wont let me put | v .5 (the point)

OpenStudy (anonymous):

a/60*a?

OpenStudy (natsent06):

what do i do?!?!!?

OpenStudy (dangerousjesse):

Yep, your exclamation made much more sense than what I was going to say haha.

OpenStudy (dangerousjesse):

Explanation*

OpenStudy (anonymous):

post a screen shot if you like

OpenStudy (natsent06):

hold on

OpenStudy (anonymous):

but in any case if it really says \[30=a\times 60\] then \[a=\frac{30}{60}=\frac{1}{2}\] making \[\frac{30}{a}=\frac{30}{\frac{1}{2}}=30\times 2=60\]

OpenStudy (anonymous):

right

OpenStudy (natsent06):

here is screenshot it only lets me put in 5

OpenStudy (natsent06):

tell me if it is not working pic or if it is there

OpenStudy (anonymous):

yeah it works

OpenStudy (natsent06):

ok good i have never taken screenshot before sooo yea lol

OpenStudy (anonymous):

\[60\]!

OpenStudy (natsent06):

what?

OpenStudy (natsent06):

um

OpenStudy (natsent06):

satellite?

OpenStudy (anonymous):

the answer is sixty \(\huge 60\)

OpenStudy (natsent06):

how

OpenStudy (dangerousjesse):

But thirty isn't equal to 120?

OpenStudy (dangerousjesse):

30=60*60 30/120

OpenStudy (natsent06):

im so confused you cant multiply 60 and 60 and get 30

OpenStudy (natsent06):

i have such a bad headache

OpenStudy (dangerousjesse):

30=0.5*60 30/0.5 a=0.5

OpenStudy (natsent06):

i know that

OpenStudy (natsent06):

but look at the screenshot it wont let me put the . it only lets me put in the 5

OpenStudy (dangerousjesse):

Can you put in 1/2?

OpenStudy (amistre64):

its not asking you to solve for 'a'; its asking you to solve for '30/a'

OpenStudy (natsent06):

no i cant put in 1/2

OpenStudy (dangerousjesse):

Ah, I thought she was replacing the variable 'a'. The answer is 60, then.

OpenStudy (natsent06):

oh ok

OpenStudy (natsent06):

wow lol

OpenStudy (natsent06):

im embarrassed

OpenStudy (amistre64):

its even stated in the original post :) suppose 30=a*60. So what is 30/a (30 over a) ^^^^^^^^^^^^^

OpenStudy (natsent06):

thanks guys sorry for the troubles

OpenStudy (amistre64):

good luck :)

OpenStudy (dangerousjesse):

Yes, I noticed haha. I merely read the question incorrectly.

OpenStudy (natsent06):

same

OpenStudy (natsent06):

sorry satellite